# Data-adaptive structural change-point detection via isolation

### v1

@@ Appendix B Proof of Theorem 1
Before proving Theorem 1, we introduce some more notation. Similar to (9), we denote the CUSUM of the unobserved signal as $\tilde{f}^{b}_{s,e}=\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum^{b}_{t=s}f_{t}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum^{e}_{t=b+1}f_{t}$. For the following proof, the contrast vector $\boldsymbol{\psi_{s,e}^{b}}=\left(\psi_{s,e}^{b}(1),\psi_{s,e}^{b}(2),\dots,\psi_{s,e}^{b}(T_{1})\right)$ is defined through the contrast function
@@
At the beginning of DAIS, $s=1,e=T$ and the first change-point that will get detected depends on the value of $d_{1,T}$. As already explained in Section 2.3, the largest difference $d_{s,e}$ will be at most at a distance $\delta_{1,T}^{j}$ from the nearest change-point, where $\delta_{1,T}^{j}\leq\frac{\delta_{j}}{2}-\frac{3\delta_{T}}{4n}$ for $\delta_{j}=|r_{j+1}-r_{j}|$. The first point to get detected will be the point that is closest to the largest difference $d_{1,T}$. The interval where the detection of this change-point occurs, cannot contain more than one change-points, as was explained in Section 2.3 and is proved at (B).
Without loss of generality, we suppose that the first change-point to get detected is $r_{J}$ for $J\in\left\{1,2,\ldots,N\right\}$. We will show that there exists an interval $[c_{m}^{l},c_{k}^{r}]$, for $m,k\in\{0,1,\ldots,K^{\max}\}$, such that $r_{J}$ is isolated assuming that $d_{1,T}>r_{J}$ and the same approach works for $d_{1,T}\leq r_{J}$. Note that since the closest change-point to the largest difference is $r_{J}$, it must hold that $d_{1,T}^{J}=d_{1,T}-r_{J}$. We are now considering the largest possible interval that may be checked before detection occurs in the sense that the largest number of expansions will be performed. Detection is guaranteed when checking this interval, but it may occur in any interval smaller than this. Showing that the change-point will be isolated in the largest interval means that it will always be isolated. Since we are considering $d_{1,T}>r_{J}$, the worst case scenario occurs for $m=k$, $c_{m}^{l}\in I^{L}_{J}$ and $c_{k}^{r}-r_{J}>\delta_{T}/2n$, but it doesn’t necessarily hold that $c_{k}^{r}\in I^{R}_{J}$. So, $r_{J}-c_{m}^{l}\leq\delta_{T}/n$ and $c_{k}^{r}-d_{1,T}+1=d_{1,T}-c_{m}^{l}$. It follows that
@@
$<2\Bigl{(}\delta_{1,T}^{J}+\frac{\delta_{T}}{n}\Bigr{)}\leq\delta_{J}+\frac{\delta_{T}}{2n}.$ (27)
Now, since $r_{J}-c_{m}^{l}\leq\delta_{T}/n$, there can be at most one change-point in the interval $[c_{m}^{l},r_{J}]$. Considering the interval $[r_{J},c_{k}^{r}]$, using (B) and $r_{J}-c_{m}^{l}>\delta_{T}/2n$, it holds that $c_{k}^{r}-r_{J}=c_{k}^{r}-c_{m}^{l}+c_{m}^{l}-r_{J}<\delta_{J}$, so $c_{k}^{r}<r_{J+1}$ and $r_{J}$ is isolated in $[r_{J},c_{k}^{r}]$. The two results combined prove that the change-point is isolated in the interval $[c_{m}^{l},c_{k}^{r}]$.
We will now show that for $\tilde{b}=\textrm{argmax}_{c_{m}^{l}\leq t<c_{k}^{r}}\lvert\tilde{X}^{t}_{c_{m}^{l},c_{k}^{r}}\rvert$, it holds that $\Big{|}\tilde{X}^{\tilde{b}}_{c_{m}^{l},c_{k}^{r}}\Big{|}>\zeta_{T}$. Using (8), we have that
@@
  • If $d_{1,T}\leq r_{J}$, then $c_{k}^{r}\in I^{R}_{J}$ and so $c_{k}^{r}-r_{J}>\delta_{T}/2n$ and $r_{J}-c_{m}^{l}+1>r_{J}-d_{1,T}+d_{1,T}-c_{m}^{l}+1\geq c_{k}^{r}-r_{J}>\delta_{T}/2n$.
@@
Therefore, we have proved that there will be an interval of the form $[c_{\tilde{m}}^{l},c_{\tilde{k}}^{r}]$, such that the interval contains $r_{J}$ and no other change-point and $\max_{c_{\tilde{m}}^{l}\leq t<c_{\tilde{k}}^{r}}\Big{|}\tilde{X}^{t}_{c_{\tilde{m}}^{l},c_{\tilde{k}}^{r}}\Big{|}>\zeta_{T}$. For $k^{\ast}\in\{0,1,\ldots K^{\max}\}$ and $m^{\ast}\in\{k^{\ast}-1,k^{\ast}\}$, denote by $c_{m^{\ast}}^{l}\geq c_{\tilde{m}}^{l}$ and $c_{k^{\ast}}^{r}\leq c_{\tilde{k}}^{r}$ the first left- and right-expanding points, respectively, that this happens and let $b_{J}=\textrm{argmax}_{c_{m^{\ast}}^{l}\leq t<c_{k^{\ast}}^{r}}\Big{|}\tilde{X}^{t}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}\Big{|}$, with $\Big{|}\tilde{X}^{b_{J}}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}\Big{|}>\zeta_{T}$. Note that $b_{J}$ cannot be an estimation of $r_{j},j\neq J$, as $r_{J}$ is isolated in the interval where it is detected. Our aim now is to find $\gamma_{T}>0$, such that for any $b^{\ast}\in\{c_{m^{\ast}}^{l},c_{m^{\ast}}^{l}+1,\ldots,c_{k^{\ast}}^{r}-1\}$ with $\lvert b^{\ast}-r_{J}\rvert\Bigl{(}\Delta_{J}^{f}\Bigl{)}^{2}>\gamma_{T}$, we have that
@@
Without loss of generality, assume that $b^{\ast}\in[r_{J},c_{k^{\ast}}^{r})$ and a similar approach holds when $b^{\ast}\in[c_{m^{\ast}}^{l},r_{J})$. Using Lemma 1, we have that for the left-hand side of (B),
@@
Using (35), (36) and (B), we can conclude that (B) is satisfied if $\Lambda>8\log T+\sqrt{2}\sqrt{\Lambda}\sqrt{8\log T}$ is satisfied, which has solution
@@
Step 4: After the detection of the change-point $r_{J}$ at the estimated location $b_{J}$ in the interval $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$, the process is repeated in the disjoint intervals $[1,c_{m^{\ast}}^{l}]$ and $[c_{k^{\ast}}^{r},T]$, which contain $r_{1},r_{2},\ldots,r_{J-1}$ and $r_{J+1},r_{J+2},\ldots,r_{N}$ respectively. This means that we do not check if there are any more change-points in the interval $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$, besides the already detected $r_{J}$. Since $c_{m^{\ast}}^{l}\geq c_{\tilde{m}}^{l}$ and $c_{k^{\ast}}^{r}\leq c_{\tilde{k}}^{r}$, it holds that $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]\subset[c_{\tilde{m}}^{l},c_{\tilde{k}}^{r}]$ and the argument of (B) can be used to conclude that $r_{J}$ is isolated in $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$.
@@
Now, we need to discuss the case that the next change-point to get detected is $r_{J-1}$. This is the closest change-point to the already detected $r_{J}$ so the end-point of the interval where DAIS is reapplied to depends on the interval where the detection of $r_{J}$ occurred. As before, for $k_{J-1},m_{J-1}\in\{0,1,\ldots,K^{\max}\}$ we will show that $r_{J-1}$ gets detected in $[c_{m_{J-1}^{\ast}}^{l},c_{k_{J-1}^{\ast}}^{r}]$, where $c_{m_{J-1}^{\ast}}^{l}\geq c_{m_{J-1}}^{l}$ and $c_{k_{J-1}^{\ast}}^{r}\leq c_{k_{J-1}}^{r}\leq c_{m^{\ast}}^{l}$ and its detection is at location
@@
which satisfies $\bigl{|}r_{J-1}-b_{J-1}\bigl{|}\Bigl{(}\Delta_{J-1}^{f}\Bigl{)}^{2}\leq C_{3}\log T$. Firstly, $r_{J-1}$ is isolated in the interval $[c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}]$ using the same argument as in (B). Following similar steps as in (29), we have that for $\tilde{b}_{J-1}=\textrm{argmax}_{c_{m_{J-1}}^{l}\leq t<c_{k_{J-1}}^{r}}\Bigl{|}\tilde{X}^{t}_{c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}}\Bigl{|}$,
@@
$c_{m^{\ast}}^{l}-r_{J-1}\geq c_{\tilde{m}}^{l}-r_{J}+\delta_{J-1}>-\frac{\delta_{T}}{n}+1+\delta_{T}>\frac{n-1}{n}\delta_{T}>\frac{\delta_{T}}{2n}$
@@
$\leq 2\Bigl{(}\frac{\delta_{J-1}}{2}-\frac{3\delta_{T}}{4n}\Bigr{)}+\frac{\delta_{T}}{n}=\delta_{J-1}-\frac{\delta_{T}}{2n}$
@@
$c_{m^{\ast}}^{l}-r_{J-1}\geq c_{\tilde{m}}^{l}-r_{J}+\delta_{J-1}\geq-\delta_{J-1}+\frac{\delta_{T}}{2n}+\delta_{J-1}=\frac{\delta_{T}}{2n}$
This means that the right end-point of the interval could eventually satisfy $c_{k_{J-1}}^{r}-r_{J-1}\geq\delta_{T}/2n$ for some $k_{J-1}$ since the largest it can become is $c_{k_{J-1}}^{r}=c_{m^{\ast}}^{l}$. As in Step 3, the detection could happen before this is satisfied. For the left end-point of the interval, the same holds as in Step 3. In any case, $\min\{c_{k_{J-1}}^{r}-r_{J-1},r_{J-1}-c_{m_{J-1}}^{l}+1\}\geq\delta_{T}/2n$ holds and so, from (B) we have that
@@
Appendix C Further simulations
Some further simulation results are presented in Tables 13 to 15 for piecewise-constant signals and Tables 16 to 18 for continuous piecewise-linear signals. The signals used are:
  • (S11)
    $small\_dist2$: sequence of length 1000 with 4 change-points at 485, 515, 900 and 930 with values between change-points 0, 1, 0, 1.5, 0. The standard deviation is $\sigma=1$.
  • (S12)
  • (S13)
  • (S14)
  • (S15)
    $wave5$: continuous piecewise-linear signal of length 1000 with 19 change-points at 50, 100, $\ldots$, 950 with changes in slope -1/16, -5/16, -5/8, 1, 5/16, 15/32, -5/8, -7/32, -3/4, 13/16, 5/16, 19/32, -1, -5/8, 23/32, 1/2, 15/16, -25/16, -5/4. The starting intercept is $f_{1}=-1$ and slope $f_{2}-f_{1}=1/32$. The standard deviation is $\sigma=0.6$.
  • (S16)
    $wave6$: continuous piecewise-linear signal of length 350 with 50 change-points at 7, 14, $\ldots$, 343 with changes in slope -2.5, 2.5, $\ldots$, 2.5, -2.5. The starting intercept is $f_{1}=-0$ and slope $f_{2}-f_{1}=1$. The standard deviation is $\sigma=1$.
The general conclusions form the tables are that DAIS performs at least as well as the competitors, often having an advantage in computational time and accuracy compared to the best performing competitors. Tables 13 and 14 indicate that in these difficult structures, where pairs of change-points are close and move in opposite directions such that they ‘cancel’ each-other out, DAIS and MOSUM have the best performance, in terms of accuracy in the number of change-points detected and MSE. For the signal (S13), in Table 15, all algorithms perform similarly, with DAIS being one of the fastest methods. The low computational complexity of DAIS can also be seen in Tables 16, 17, 18.
Table 13: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S11). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $\leq-3$ $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 17 2 78 3 0 0 0.0220 0.084 0.004
ID_th 0 3 1 69 10 11 6 0.0228 0.071 0.004
ID_ic 0 26 0 71 3 0 0 0.0230 0.114 0.002
WBS_th 0 0 0 17 18 18 47 0.0349 0.167 0.022
WBS_ic 0 27 0 71 2 0 0 0.0221 0.118 0.022
WBS2 0 6 3 70 8 7 6 0.0221 0.075 0.485
PELT 9 72 0 19 0 0 0 0.0389 0.378 0.001
NOT 0 25 0 72 2 1 0 0.0217 0.111 0.024
MOSUM 0 8 1 85 4 2 0 0.0195 0.055 0.012
MSCP 89 9 1 1 0 0 0 0.0836 0.576 158.370
Table 14: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S12). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $\leq-3$ $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 15 3 80 0 2 0 0.0321 0.067 0.004
ID_th 0 2 2 75 8 7 6 0.0323 0.037 0.003
ID_ic 2 34 1 62 1 0 0 0.0350 0.143 0.002
WBS_th 0 0 0 22 24 19 35 0.0414 0.092 0.022
WBS_ic 2 31 1 65 0 1 0 0.0331 0.132 0.022
WBS2 0 6 5 70 9 6 4 0.0306 0.132 0.479
PELT 26 65 0 9 0 0 0 0.0596 0.437 0.001
NOT 2 30 1 67 0 0 0 0.0333 0.128 0.024
MOSUM 0 13 2 79 3 3 0 0.0287 0.063 0.015
MSCP 98 2 0 0 0 0 0 0.1457 0.632 158.334
Table 15: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S13). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $\leq-3$ $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 0 1 94 4 1 0 2.72 ×$\times 10^{-2}$ 0.010 0.002
ID_th 0 0 1 84 12 2 1 2.72 ×$\times 10^{-2}$ 0.011 0.001
ID_ic 0 0 0 87 12 1 0 2.46 ×$\times 10^{-2}$ 0.009 0.002
WBS_th 0 0 0 58 16 19 7 2.58 ×$\times 10^{-2}$ 0.015 0.014
WBS_ic 0 0 0 92 8 0 0 2.35 ×$\times 10^{-2}$ 0.008 0.014
WBS2 0 0 1 96 3 0 0 2.31 ×$\times 10^{-2}$ 0.008 0.088
PELT 1 6 3 90 0 0 0 2.63 ×$\times 10^{-2}$ 0.016 0.001
NOT 0 0 0 95 5 0 0 2.47 ×$\times 10^{-2}$ 0.009 0.021
MOSUM 0 0 0 95 3 2 0 2.26 ×$\times 10^{-2}$ 0.008 0.008
MSCP 100 0 0 0 0 0 0 22.49 ×$\times 10^{-2}$ 0.295 14.850
Table 16: Distribution of $\hat{N}-N$ of 100 simulated data sequences of (S14). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $\leq-15$ $(-15,-2]$ $-1$ 0 1 $[2,15)$ $\geq 15$ MSE $d_{H}$ Time (s)
DAIS 0 0 0 96 3 1 0 0.025 0.116 0.002
ID_th 0 0 0 65 23 12 0 0.020 0.188 0.001
ID_ic 0 0 0 99 1 0 0 0.013 0.108 0.003
CPOP 0 0 0 90 6 4 0 0.083 0.125 0.014
NOT 0 0 1 95 2 2 0 0.014 0.102 0.070
MARS 0 91 7 2 0 0 0 3.990 2.260 0.003
TF 0 0 0 0 0 35 65 0.026 0.432 0.082
TS 0 0 0 99 0 1 0 0.093 0.195 0.079
Table 17: Distribution of $\hat{N}-N$ of 100 simulated data sequences of (S15). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $\leq-15$ $(-15,-2]$ $-1$ 0 1 $[2,15)$ $\geq 15$ MSE $d_{H}$ Time (s)
DAIS 0 0 4 84 12 0 0 0.057 0.147 0.006
ID_th 0 0 0 55 38 7 0 0.059 0.150 0.047
ID_ic 0 0 0 87 13 0 0 0.038 0.108 0.011
CPOP 0 0 0 90 10 0 0 0.089 0.086 0.187
NOT 0 0 60 37 3 0 0 0.043 0.623 0.238
MARS 0 100 0 0 0 0 0 22.000 1.610 0.005
TF 0 0 0 0 0 0 100 0.032 0.467 0.511
TS 0 0 7 85 7 1 0 0.225 0.225 0.468
Table 18: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S16). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $\leq-15$ $(-15,-2]$ $-1$ 0 1 $[2,15)$ $\geq 15$ MSE $d_{H}$ Time (s)
DAIS 0 1 5 94 0 0 0 0.497 0.304 0.007
ID_th 0 0 2 91 7 0 0 0.496 0.290 0.004
ID_ic 37 3 9 51 0 0 0 3.000 18.500 0.022
CPOP 0 0 0 96 3 1 0 1.660 0.316 0.023
NOT 100 0 0 0 0 0 0 6.650 49.000 0.036
MARS 100 0 0 0 0 0 0 6.590 47.800 0.003
TF 0 0 0 0 0 72 28 1.210 0.319 0.102
TS 100 0 0 0 0 0 0 6.650 49.000 0.157

### v2

@@ Appendix C Proof of Theorem 1
Before proving Theorem 1, we introduce some more notation. For convinience, we denote (10) by $\tilde{X}^{b}_{s,e}=\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum^{b}_{t=s}X_{t}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum^{e}_{t=b+1}X_{t}.$ Similarly, we denote the CUSUM of the unobserved signal as $\tilde{f}^{b}_{s,e}=\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum^{b}_{t=s}f_{t}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum^{e}_{t=b+1}f_{t}$. For the following proof, the contrast vector $\boldsymbol{\psi_{s,e}^{b}}=\left(\psi_{s,e}^{b}(1),\psi_{s,e}^{b}(2),\dots,\psi_{s,e}^{b}(T_{1})\right)$ is defined through the contrast function
@@
At the beginning of DAIS, $s=1,e=T$ and the first change-point that will get detected depends on the value of $d_{1,T}$. As already explained in Section 2.3, for $j\in\{1,\ldots,N-1\}$, the largest difference $d_{s,e}$ will be at most at a distance $\delta_{1,T}^{j}$ from the nearest change-point $r_{j}$ or $r_{j+1}$, where $\delta_{1,T}^{j}\leq\frac{\tilde{\delta}_{j}}{2}-\frac{3\delta_{T}}{4n}$ for $\tilde{\delta}_{j}=r_{j+1}-r_{j}$. The first point to get detected will be the point that is closest to the largest difference $d_{1,T}$. The interval where the detection of this change-point occurs, cannot contain more than one change-points, as was explained in Section 2.3 and is proved at (C).
We will show that there exists an interval $[c_{m}^{l},c_{k}^{r}]$, for $m,k\in\{0,1,\ldots,K^{\max}\}$, such that the first change-point to get detected, $r_{J}$, is isolated, assuming that $d_{1,T}>r_{J}$. The same approach works for $d_{1,T}\leq r_{J}$. Without loss of generality, we suppose that $J\in\left\{1,2,\ldots,N-1\right\}$. The result in the case that $J=N$ is easier to prove and so it is skipped. Note that since the closest change-point to the largest difference is $r_{J}$, it must hold that $d_{1,T}^{J}=d_{1,T}-r_{J}$. We are now considering the largest possible interval that may be checked before detection occurs in the sense that the largest number of expansions will be performed. Detection is guaranteed when checking this interval, but it may occur in any interval smaller than this. Showing that the change-point will be isolated in the largest interval means that it will always be isolated. Since we are considering $d_{1,T}>r_{J}$, the worst case scenario occurs for $m=k$, $c_{m}^{l}\in I^{L}_{J}$ and $c_{k}^{r}-r_{J}>\delta_{T}/2n$, but it doesn’t necessarily hold that $c_{k}^{r}\in I^{R}_{J}$. So, $r_{J}-c_{m}^{l}\leq\delta_{T}/n$ and $c_{k}^{r}-d_{1,T}+1=d_{1,T}-c_{m}^{l}$. It follows that
@@
$<2\Bigl{(}\delta_{1,T}^{J}+\frac{\delta_{T}}{n}\Bigr{)}\leq\tilde{\delta}_{J}+\frac{\delta_{T}}{2n}.$ (33)
Now, since $r_{J}-c_{m}^{l}\leq\delta_{T}/n$, there can be at most one change-point in the interval $[c_{m}^{l},r_{J}]$. Considering the interval $[r_{J},c_{k}^{r}]$, using (C) and $r_{J}-c_{m}^{l}>\delta_{T}/2n$, it holds that $c_{k}^{r}-r_{J}=c_{k}^{r}-c_{m}^{l}+c_{m}^{l}-r_{J}<\tilde{\delta}_{J}$, so $c_{k}^{r}<r_{J+1}$ and $r_{J}$ is isolated in $[r_{J},c_{k}^{r}]$. The two results combined prove that the change-point is isolated in the interval $[c_{m}^{l},c_{k}^{r}]$.
Now, we consider $J\in\{1,\ldots,N\}$. We will show that for $\tilde{b}=\textrm{argmax}_{c_{m}^{l}\leq t<c_{k}^{r}}\lvert\tilde{X}^{t}_{c_{m}^{l},c_{k}^{r}}\rvert$, it holds that $\Big{|}\tilde{X}^{\tilde{b}}_{c_{m}^{l},c_{k}^{r}}\Big{|}>\zeta_{T}$. Using (29), we have that
@@
  • If $d_{1,T}\leq r_{J}$, then $c_{k}^{r}\in I^{R}_{J}$ and so $c_{k}^{r}-r_{J}>\delta_{T}/2n$ and $r_{J}-c_{m}^{l}+1=r_{J}-d_{1,T}+d_{1,T}-c_{m}^{l}+1>c_{k}^{r}-d_{1,T}\geq c_{k}^{r}-r_{J}>\delta_{T}/2n$.
@@
Therefore, we have proved that there will be an interval of the form $[c_{\tilde{m}}^{l},c_{\tilde{k}}^{r}]$, such that the interval contains $r_{J}$ and no other change-point and $\max_{c_{\tilde{m}}^{l}\leq t<c_{\tilde{k}}^{r}}\Big{|}\tilde{X}^{t}_{c_{\tilde{m}}^{l},c_{\tilde{k}}^{r}}\Big{|}>\zeta_{T}$. For $k^{\ast}\in\{1,2,\ldots K^{\max}\}$ and $m^{\ast}\in\{k^{\ast}-1,k^{\ast}\}$, denote by $c_{m^{\ast}}^{l}\geq c_{\tilde{m}}^{l}$ and $c_{k^{\ast}}^{r}\leq c_{\tilde{k}}^{r}$ the first left- and right-expanding points, respectively, that this happens and let $b_{J}=\textrm{argmax}_{c_{m^{\ast}}^{l}\leq t<c_{k^{\ast}}^{r}}\Big{|}\tilde{X}^{t}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}\Big{|}$, with $\Big{|}\tilde{X}^{b_{J}}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}\Big{|}>\zeta_{T}$. Note that $b_{J}$ cannot be an estimation of $r_{j},j\neq J$, as $r_{J}$ is isolated in the interval where it is detected. Our aim now is to find $\gamma_{T}>0$, such that for any $b^{\ast}\in\{c_{m^{\ast}}^{l},c_{m^{\ast}}^{l}+1,\ldots,c_{k^{\ast}}^{r}-1\}$ with $\lvert b^{\ast}-r_{J}\rvert\Bigl{(}\Delta_{J}^{f}\Bigl{)}^{2}>\gamma_{T}$, we have that
@@
Without loss of generality, assume that $b^{\ast}\in[r_{J},c_{k^{\ast}}^{r})$ and a similar approach holds when $b^{\ast}\in[c_{m^{\ast}}^{l},r_{J})$. Using Lemma 1, we have that for the left-hand side of (C),
@@
Using (41), (42) and (C), we can conclude that (C) is satisfied if $\Lambda>8\log T+\sqrt{2}\sqrt{\Lambda}\sqrt{8\log T}$ is satisfied, which has solution
@@
Step 4: After the detection of the change-point $r_{J}$ at the estimated location $b_{J}$ in the interval $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$, the process is repeated in the disjoint intervals $[1,c_{m^{\ast}}^{l}]$ and $[c_{k^{\ast}}^{r},T]$, which contain $r_{1},r_{2},\ldots,r_{J-1}$ and $r_{J+1},r_{J+2},\ldots,r_{N}$ respectively. This means that we do not check if there are any more change-points in the interval $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$, besides the already detected $r_{J}$. Since $c_{m^{\ast}}^{l}\geq c_{\tilde{m}}^{l}$ and $c_{k^{\ast}}^{r}\leq c_{\tilde{k}}^{r}$, it holds that $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]\subset[c_{\tilde{m}}^{l},c_{\tilde{k}}^{r}]$ and the argument of (C) can be used to conclude that $r_{J}$ is isolated in $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$.
@@
Now, we need to discuss the case that the next change-point to get detected is $r_{J-1}$. This is the closest change-point to the already detected $r_{J}$ so the end-point of the interval where DAIS is reapplied to depends on the interval where the detection of $r_{J}$ occurred. As before, for $k_{J-1}\in\{1,\ldots,K^{\max}\}$ and $m_{J-1}\in\{k_{J-1}-1,k_{J-1}\}$ we will show that $r_{J-1}$ gets detected in $[c_{m_{J-1}^{\ast}}^{l},c_{k_{J-1}^{\ast}}^{r}]$, where $c_{m_{J-1}^{\ast}}^{l}\geq c_{m_{J-1}}^{l}$ and $c_{k_{J-1}^{\ast}}^{r}\leq c_{k_{J-1}}^{r}\leq c_{m^{\ast}}^{l}$ and its detection is at location
@@
which satisfies $\bigl{|}r_{J-1}-b_{J-1}\bigl{|}\Bigl{(}\Delta_{J-1}^{f}\Bigl{)}^{2}\leq C_{3}\log T$. Firstly, $r_{J-1}$ is isolated in the interval $[c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}]$ using the same argument as in (C). Following similar steps as in (35), we have that for $\tilde{b}_{J-1}=\textrm{argmax}_{c_{m_{J-1}}^{l}\leq t<c_{k_{J-1}}^{r}}\Bigl{|}\tilde{X}^{t}_{c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}}\Bigl{|}$,
@@
$c_{m^{\ast}}^{l}-r_{J-1}\geq c_{\tilde{m}}^{l}-r_{J}+\tilde{\delta}_{J-1}>-\frac{\delta_{T}}{n}+1+\delta_{T}>\frac{n-1}{n}\delta_{T}>\frac{\delta_{T}}{2n}$
@@
$\leq 2\Bigl{(}\frac{\tilde{\delta}_{J-1}}{2}-\frac{3\delta_{T}}{4n}\Bigr{)}+\frac{\delta_{T}}{n}=\tilde{\delta}_{J-1}-\frac{\delta_{T}}{2n}$
@@
$c_{m^{\ast}}^{l}-r_{J-1}\geq c_{\tilde{m}}^{l}-r_{J}+\tilde{\delta}_{J-1}\geq-\tilde{\delta}_{J-1}+\frac{\delta_{T}}{2n}+\tilde{\delta}_{J-1}=\frac{\delta_{T}}{2n}$
This means that the right end-point of the interval could eventually satisfy $c_{k_{J-1}}^{r}-r_{J-1}\geq\delta_{T}/2n$ for some $k_{J-1}$ since the largest it can become is $c_{k_{J-1}}^{r}=c_{m^{\ast}}^{l}$. As in Step 3, the detection could happen before this is satisfied. For the left end-point of the interval, the same holds as in Step 3. In any case, $\min\{c_{k_{J-1}}^{r}-r_{J-1},r_{J-1}-c_{m_{J-1}}^{l}+1\}\geq\delta_{T}/2n$ holds and so, from (C) we have that